OpenStudy (anonymous):

Factor completely: 2x3 + 10x2 + 4x + 20

OpenStudy (nurali):

2x^3 + 10x^2 + 4x + 20 = 0 (2x+10)(x^2+2) = 0 2x = -10 ; x^2 = -2 x = -5 ; x = sqrt of -2 x = -5 ; x = + or - i sqrt of 2

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

u have the ans

OpenStudy (nurali):

\[x=-5 \] \[x=+-i \sqrt{2}\]

OpenStudy (anonymous):

im not understanding what the answer is can u pls break it down for me

OpenStudy (anonymous):

did u understand the procedure

OpenStudy (anonymous):

not at all i just need the answer

OpenStudy (anonymous):

with understanding how will u get the ans .... u will never get ur ans ,.. understand the procedure first ... ur ans is given.....